OpenStudy (anonymous):

Estimate the volume of juice in ounces that the cylinder will hold. Round to the nearest tenth of an ounce. (Hint: 1 in3 ≈ 0.55 oz.)

OpenStudy (aum):

Do they tell you the length of the straw somewhere?

OpenStudy (aum):

Volume of the cylinder \(\Large V = \pi r^2h\) \(\Large r = 2\). We need h. We can find h if we know the length of the straw.

OpenStudy (anonymous):

the length of the straw is 6 inches in total.

OpenStudy (anonymous):

and i tried to plug it all in the equation there was no answer like that

OpenStudy (aum):



OpenStudy (aum):

h^2 + 4^2 = 5^2 h^2 + 16 = 25 h^2 = 25 - 16 = 9 h = 3 V = pi * r^2 * h = pi * 2^2 * 3 = 12pi

OpenStudy (anonymous):

h=3

OpenStudy (aum):

V = 12 * pi = 37.70 in^3 1 in^3 = 0.55 oz 37.70 in^3 = 37.7 * 0.55 = 20.7 oz.

OpenStudy (anonymous):

ohh ok thanks thas what i just got right now too!!

OpenStudy (aum):

Cool. You are welcome.

OpenStudy (anonymous):

do you mind helping me on one more plz?

OpenStudy (aum):

I can try one more.

OpenStudy (anonymous):

Estimate the volume of juice in ounces that the prism will hold. Round to the nearest tenth of an ounce. (Hint: 1 in3 ≈ 0.55 oz.)

OpenStudy (anonymous):

im thinking that the straw length would also be the same

OpenStudy (aum):

Find the height of the prism using the same method we used before. Then, volume of the prism = 3 x 3 x h in^3 Use the conversion factor of 0.55 to convert in^3 to oz.

OpenStudy (aum):

Finding the height is a bit trickier than before. Find the diagonal length of the bottom square first.

OpenStudy (aum):



OpenStudy (anonymous):

i got 19.8 but it also want s me to round

OpenStudy (anonymous):

to the nearest tenth

OpenStudy (aum):

What was the height of the prism that you got?

OpenStudy (anonymous):

9

OpenStudy (aum):

That may not be correct. In my diagram above, AC^2 = 3^2 + 3^2 = 9+9 = 18 Then for the prism, AC^2 + h^2 = 5^2 (assume the length of the straw inside the prism is 5 in.) So 18 + h^2 = 25. h^2 = 25 - 18 = 7. h = sqrt(7). Volume of prism = 3 x 3 x sqrt(7) = 23.81 in^3 1 in^3 = 0.55 oz 23.81 in^3 = 23.81 * 0.55 = 13.1 oz.

OpenStudy (anonymous):

DUDE ITS LIKE EINSTEIN WAS REBORN AGAIN!! UR A GENIUS THANKS MAN!!!

OpenStudy (aum):

lol. You are welcome.